\begin{thebibliography}{99}


\bibitem{Adler}
  S.L.~Adler,
  % ``Some simple vacuum--polarization phenomenology: $\epem\!\!\to\:$hadrons;
  % the muonic--atom $X$--ray discrepancy and ($g_{\mu}-2$)'',
  Phys.\ Rev.\ D~\textbf{10}, 3714 (1974).
  % doi:10.1103/PhysRevD.10.3714
  %%CITATION = doi:10.1103/PhysRevD.10.3714;%%


\bibitem{Rad82}
  A.V.~Radyushkin,
  % ``Optimized $\Lambda$--parametrization for the QCD running
  % coupling constant in spacelike and timelike regions'',
  report JINR E2--82--159 (1982); JINR Rapid Commun.~\textbf{78}, 96 (1996);
  arXiv:hep-ph/9907228.
  %%CITATION = HEP-PH/9907228;%%


\bibitem{KP82}
  N.V.~Krasnikov and A.A.~Pivovarov,
  % ``The influence of the analytic continuation effects on the value of the QCD scale
  % parameter $\Lambda$ extracted from charmonium and upsilonium hadron decays'',
  Phys.\ Lett.\ B~\textbf{116}, 168 (1982).
  % doi:10.1016/0370-2693(82)91001-2
  %%CITATION = doi:10.1016/0370-2693(82)91001-2;%%


\bibitem{Book}
  A.V.~Nesterenko, \textit{``Strong interactions in spacelike and timelike
  domains: Dispersive approach''} (Elsevier, Amsterdam, 222~p.,~2017).
  %%CITATION = INSPIRE-1514185;%%


\bibitem{DPT1}
  A.V.~Nesterenko and J.~Papavassiliou,
  % ``A novel integral representation for the Adler function'',
  J.~Phys.\ G~\textbf{32}, 1025 (2006).
  % doi:10.1088/0954-3899/32/7/011
  % arXiv:hep-ph/0511215.
  %%CITATION = doi:10.1088/0954-3899/32/7/011;%%


\bibitem{DPT2}
  A.V.~Nesterenko,
  % ``Dispersive approach to QCD and inclusive tau lepton hadronic decay'',
  Phys.\ Rev.\ D~\textbf{88}, 056009 (2013);
  % doi:10.1103/PhysRevD.88.056009
  % arXiv:1306.4970~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevD.88.056009;%%
%
  % A.V.~Nesterenko,
  % ``Hadronic vacuum polarization function within dispersive approach to QCD'',
  J.~Phys.\ G~\textbf{42}, 085004 (2015).
  % doi:10.1088/0954-3899/42/8/085004
  % arXiv:1411.2554~[hep-ph].
  %%CITATION = doi:10.1088/0954-3899/42/8/085004;%%


\bibitem{DPTPrelim}
  A.V.~Nesterenko and J.~Papavassiliou,
  % ``The massive analytic invariant charge in QCD'',
  Phys.\ Rev.\ D~\textbf{71}, 016009 (2005);
  % doi:10.1103/PhysRevD.71.016009
  % arXiv:hep-ph/0410406.
  %%CITATION = doi:10.1103/PhysRevD.71.016009;%%
%
  % A.V.~Nesterenko and J.~Papavassiliou,
  % ``The QCD analytic effective charge and its dependence on the pion mass'',
  Int.\ J.\ Mod.\ Phys.\ A~\textbf{20}, 4622 (2005);
  % doi:10.1142/S0217751X05028296
  % arXiv:hep-ph/0409220.
  %%CITATION = doi:10.1142/S0217751X05028296;%%
%
  % A.V.~Nesterenko and J.~Papavassiliou,
  % ``The QCD analytic running coupling and chiral symmetry breaking'',
  Nucl.\ Phys.\ B (Proc.\ Suppl.) \textbf{152}, 47 (2005);
  % doi:10.1016/j.nuclphysbps.2005.08.009
  % arXiv:hep-ph/0410072.
  %%CITATION = doi:10.1016/j.nuclphysbps.2005.08.009;%%
%
  % A.V.~Nesterenko and J.~Papavassiliou,
  % ``Impact of the pion mass on nonpower expansion for QCD observables'',
  % Nucl.\ Phys.\ B (Proc.\ Suppl.)
  \textbf{164}, 304 (2007).
  % doi:10.1016/j.nuclphysbps.2006.11.089
  % arXiv:hep-ph/0507320.
  %%CITATION = doi:10.1016/j.nuclphysbps.2006.11.089;%%


\bibitem{DPT3}
  A.V.~Nesterenko,
  % ``On the low--energy behavior of the Adler function'',
  Nucl.\ Phys.\ B (Proc.\ Suppl.) \textbf{186}, 207 (2009);
  % doi:10.1016/j.nuclphysbps.2008.12.048
  % arXiv:0808.2043~[hep-ph].
  %%CITATION = doi:10.1016/j.nuclphysbps.2008.12.048;%%
%
  % A.V.~Nesterenko,
  % ``Hadronic effects in low--energy QCD: inclusive tau lepton decay'',
  % Nucl.\ Phys.\ B (Proc.\ Suppl.)
  \textbf{234}, 199 (2013);
  % doi:10.1016/j.nuclphysbps.2012.12.013
  % arXiv:1209.0164~[hep-ph].
  %%CITATION = doi:10.1016/j.nuclphysbps.2012.12.013;%%
%
  % A.V.~Nesterenko,
  % ``Dispersive approach to QCD: $\tau$~lepton hadronic decay in vector
  % and axial--vector channels'',
  Nucl.\ Part.\ Phys.\ Proc.~\textbf{258}, 177 (2015);
  % doi:10.1016/j.nuclphysbps.2015.01.038
  % arXiv:1409.0687~[hep-ph].
  %%CITATION = doi:10.1016/j.nuclphysbps.2015.01.038;%%
%
  % A.V.~Nesterenko,
  % ``Hadronic contributions to electroweak observables within DPT'',
  % Nucl.\ Part.\ Phys.\ Proc.~
  \textbf{270}, 206 (2016);
  % doi:10.1016/j.nuclphysbps.2016.02.041
  % arXiv:1509.00197~[hep-ph].
  %%CITATION = doi:10.1016/j.nuclphysbps.2016.02.041;%%
%
%
  % A.V.~Nesterenko,
  % ``Adler function in the analytic approach to QCD'',
  SLAC eConf C0706044, 25 (2008);
  % arXiv:0710.5878~[hep-ph].
  %%CITATION = ARXIV:0710.5878;%%
%
  % A.V.~Nesterenko,
  % ``The effects due to hadronization in the inclusive tau lepton decay'',
  % SLAC eConf
  C1106064, 23 (2011);
  % arXiv:1106.4006~[hep-ph].
  %%CITATION = ARXIV:1106.4006;%%
%
%
  % A.V.~Nesterenko,
  % ``Inclusive tau lepton decay: the effects due to hadronization'',
  PoS~ConfinementX, 350 (2012);
  % arXiv:1302.0518~[hep-ph].
  %%CITATION = ARXIV:1302.0518;%%
%
  % A.V.~Nesterenko,
  % ``Inclusive $\tau$~lepton hadronic decay in vector and axial--vector channels
  % within dispersive approach to~QCD'',
  AIP Conf.\ Proc.~\textbf{1701}, 040016 (2016);
  % doi:10.1063/1.4938633
  % arXiv:1508.03705~[hep-ph].
  %%CITATION = doi:10.1063/1.4938633;%%
%
  % A.V.~Nesterenko,
  % ``Dispersive approach to QCD and hadronic contributions
  % to electroweak observables'',
  EPJ Web Conf.~\textbf{137}, 05021 (2017).
  % doi:10.1051/epjconf/201713705021
  % arXiv:1701.00678~[hep-ph].
  %%CITATION = doi:10.1051/epjconf/201713705021;%%


\bibitem{APT0}
  D.V.~Shirkov and I.L.~Solovtsov,
  % ``Analytic model for the QCD running coupling with universal $\alpha_{s}(0)$ value'',
  Phys.\ Rev.\ Lett.~\textbf{79}, 1209 (1997);
  % doi:10.1103/PhysRevLett.79.1209
  % arXiv:hep-ph/9704333.
  %%CITATION = doi:10.1103/PhysRevLett.79.1209;%%
%
  % D.V.~Shirkov and I.L.~Solovtsov,
  % ``Ten years of the Analytic Perturbation Theory in QCD'',
  Theor.\ Math.\ Phys.~\textbf{150}, 132 (2007);
  % doi:10.1007/s11232-007-0010-7
  % arXiv:hep-ph/0611229.
  %%CITATION = doi:10.1007/s11232-007-0010-7;%%
%
  K.A.~Milton and I.L.~Solovtsov,
  % ``Analytic perturbation theory in QCD and Schwinger's connection between
  % the beta function and the spectral density'',
  Phys.\ Rev.\ D~\textbf{55}, 5295 (1997);
  % doi:10.1103/PhysRevD.55.5295
  % arXiv:hep-ph/9611438.
  %%CITATION = doi:10.1103/PhysRevD.55.5295;%%
%
  % K.A.~Milton and I.L.~Solovtsov,
  % ``Can the QCD effective charge be symmetrical in the Euclidean
  % and the Minkowskian regions?'',
  Phys.\ Rev.\ D~\textbf{59}, 107701 (1999).
  % doi:10.1103/PhysRevD.59.107701
  % arXiv:hep-th/9812171.
  %%CITATION = doi:10.1103/PhysRevD.59.107701;%%


\bibitem{APT1}
  G.~Cvetic and C.~Valenzuela,
  % ``Various versions of analytic QCD and skeleton-motivated evaluation of observables'',
  Phys.\ Rev.\ D~\textbf{74}, 114030 (2006);
  \textbf{84}, 019902(E) (2011);
  % [Erratum-ibid.\ D {\bf 84}, 019902 (2011)]
  % arXiv:hep-ph/0608256.
  %%CITATION = HEP-PH/0608256;%%
%
  % G.~Cvetic and C.~Valenzuela,
  % ``Analytic QCD: A Short review'',
  Braz.\ J.\ Phys.~\textbf{38}, 371 (2008);
  % arXiv:0804.0872~[hep-ph].
  %%CITATION = ARXIV:0804.0872;%%
%
  G.~Cvetic and A.V.~Kotikov,
  % ``Analogs of Noninteger Powers in General Analytic QCD,''
  J.~Phys.\ G~\textbf{39}, 065005 (2012);
  % arXiv:1106.4275~[hep-ph].
  %%CITATION = ARXIV:1106.4275;%%
%
  A.P.~Bakulev,
  % ``Global Fractional Analytic Perturbation Theory in QCD
  % with Selected Applications,''
  Phys.\ Part.\ Nucl.~\textbf{40}, 715 (2009);
  % arXiv:0805.0829~[hep-ph].
  %%CITATION = ARXIV:0805.0829;%%
%
  N.G.~Stefanis,
  % ``Taming Landau singularities in QCD perturbation theory:
  % The Analytic approach'',
  % Phys.\ Part.\ Nucl.
  \textit{ibid.}~\textbf{44}, 494 (2013).
  % arXiv:0902.4805~[hep-ph].
  %%CITATION = ARXIV:0902.4805;%%


\bibitem{APT2}
  G.~Cvetic, A.Y.~Illarionov, B.A.~Kniehl, and A.V.~Kotikov,
  % ``Small-x behavior of the structure function F(2) and
  % its slope partial ln F(2) / partial ln(1/x) for 'frozen'
  % and analytic strong-coupling constants'',
  Phys.\ Lett.\ B~\textbf{679}, 350 (2009);
  % arXiv:0906.1925~[hep-ph].
  %%CITATION = ARXIV:0906.1925;%%
%
  A.V.~Kotikov,
  % ``Small $x$ behavior of parton distributions. Analytical
  % and 'frozen' coupling constants. BFKL corrections'',
  PoS~(Baldin ISHEPP~XXI), 033 (2013);
  % arXiv:1212.3733~[hep-ph].
  %%CITATION = ARXIV:1212.3733;%%
%
  % A.~Kotikov,
  % ``Parton distributions at low $x$ and gluon- and quark
  % average multiplicities,''
  PoS~(Baldin ISHEPP~XXII), 028 (2015);
  % arXiv:1502.07108~[hep-ph].
  %%CITATION = ARXIV:1502.07108;%%
%
  A.V.~Kotikov and B.G.~Shaikhatdenov,
  % ``Q2-evolution of parton densities at small x values.
  % Combined H1 and ZEUS F2 data'',
  Phys.\ Part.\ Nucl.~\textbf{44}, 543 (2013);
  % arXiv:1212.4582~[hep-ph].
  %%CITATION = ARXIV:1212.4582;%%
%
  % A.V.~Kotikov and B.G.~Shaikhatdenov,
  % ``$Q^2$--evolution of parton distributions at small values of $x$:
  % Effective scale for combined H1 and ZEUS data on the structure function $F_2$'',
  Phys.\ Atom.\ Nucl.~\textbf{78}, 525 (2015);
  % arXiv:1402.4349~[hep-ph].
  %%CITATION = ARXIV:1402.4349;%%
%
  A.V.~Kotikov, V.G.~Krivokhizhin, and B.G.~Shaikhatdenov,
  % ``Improved nonsinglet QCD analysis of fixed--target DIS data'',
  J.~Phys.\ G~\textbf{42}, 095004 (2015);
  % arXiv:1411.1236~[hep-ph].
  %%CITATION = ARXIV:1411.1236;%%
%
  A.V.~Kotikov, B.G.~Shaikhatdenov, and P.~Zhang,
  % ``Application of the rescaling model at small Bjorken $x$ values'',
  Phys.\ Rev.\ D~\textbf{96}, 114002 (2017);
  % arXiv:1706.01849~[hep-ph];
  %%CITATION = ARXIV:1706.01849;%%
%
  C.~Ayala, G.~Cvetic, A.V.~Kotikov, and B.G.~Shaikhatdenov,
  % ``Bjorken sum rule in QCD frameworks with analytic (holomorphic) coupling'',
  arXiv:1708.06284~[hep-ph].
  %%CITATION = ARXIV:1708.06284;%%


\bibitem{APT3}
  G.~Cvetic and C.~Villavicencio,
  % ``Operator Product Expansion with analytic QCD in tau decay physics'',
  Phys.\ Rev.\ D~\textbf{86}, 116001 (2012);
  % arXiv:1209.2953~[hep-ph].
  %%CITATION = ARXIV:1209.2953;%%
%
  C.~Ayala and G.~Cvetic,
  % ``Calculation of binding energies and masses of quarkonia in analytic QCD models'',
  % Phys.\ Rev.\ D
  \textit{ibid.}~\textbf{87}, 054008 (2013);
  % arXiv:1210.6117~[hep-ph].
  %%CITATION = ARXIV:1210.6117;%%
%
  P.~Allendes, C.~Ayala, and G.~Cvetic,
  % ``Gluon Propagator in Fractional Analytic Perturbation Theory'',
  % Phys.\ Rev.\ D
  \textit{ibid.}~\textbf{89}, 054016 (2014);
  % arXiv:1401.1192~[hep-ph].
  %%CITATION = ARXIV:1401.1192;%%
%
  C.~Ayala, G.~Cvetic, and R.~Kogerler,
  % ``Lattice-motivated holomorphic nearly perturbative QCD'',
  J.~Phys.\ G~\textbf{44}, 075001 (2017);
  % doi:10.1088/1361-6471/aa6fdf
  % arXiv:1608.08240~[hep-ph].
  %%CITATION = doi:10.1088/1361-6471/aa6fdf;%%
%
  C.~Ayala, G.~Cvetic, R.~Kogerler, and I.~Kondrashuk,
  % ``Nearly perturbative lattice--motivated QCD coupling with zero IR limit'',
  arXiv:1703.01321~[hep-ph];
  %%CITATION = ARXIV:1703.01321;%%
%
  C.~Ayala,
  % ``QCD coupling which respects lattice restrictions at low energies'',
  arXiv:1710.00255~[hep-ph];
  %%CITATION = ARXIV:1710.00255;%%
%
  G.~Cvetic,
  % ``Nearly perturbative QCD coupling with lattice-motivated zero IR limit'',
  arXiv:1712.00622~[hep-ph].
  %%CITATION = ARXIV:1712.00622;%%


\bibitem{APT4}
  M.~Baldicchi and G.M.~Prosperi,
  % ``Infrared behavior of the running coupling constant
  % and bound states in QCD,''
  Phys.\ Rev.\ D~\textbf{66}, 074008 (2002);
  % [hep-ph/0202172].
  %%CITATION = HEP-PH/0202172;%%
%
  % M.~Baldicchi and G.M.~Prosperi,
  % ``Running coupling constant and masses in QCD,
  % the meson spectrum,''
  AIP Conf.\ Proc.~\textbf{756}, 152 (2005);
  % [hep-ph/0412359].
  %%CITATION = HEP-PH/0412359;%%
%
  M.~Baldicchi, G.M.~Prosperi, and C.~Simolo,
  % ``Extracting infrared QCD coupling from meson spectrum,''
  % AIP Conf.\ Proc.
  \textit{ibid.}~\textbf{892}, 340 (2007);
  % [hep-ph/0611087].
  %%CITATION = HEP-PH/0611087;%%
%
  M.~Baldicchi, A.V.~Nesterenko, G.M.~Prosperi, D.V.~Shirkov, and C.~Simolo,
  % ``Bound state approach to the QCD coupling at low energy scales'',
  Phys.\ Rev.\ Lett.~\textbf{99}, 242001 (2007);
  % doi:10.1103/PhysRevLett.99.242001
  % arXiv:0705.0329~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevLett.99.242001;%%
%
  M.~Baldicchi, A.V.~Nesterenko, G.M.~Prosperi, and C.~Simolo,
  % ``QCD coupling below 1 GeV from quarkonium spectrum'',
  Phys.\ Rev.\ D~\textbf{77}, 034013 (2008).
  % doi:10.1103/PhysRevD.77.034013
  % arXiv:0705.1695~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevD.77.034013;%%


\bibitem{APT5}
  N.~Christiansen, M.~Haas, J.M.~Pawlowski, and N.~Strodthoff,
  % ``Transport coefficients in Yang--Mills theory and QCD'',
  Phys.\ Rev.\ Lett.~\textbf{115}, 112002 (2015);
  % doi:10.1103/PhysRevLett.115.112002
  % arXiv:1411.7986~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevLett.115.112002;%%
%
  N.~Mueller and J.M.~Pawlowski,
  % ``Magnetic catalysis and inverse magnetic catalysis in QCD'',
  Phys.\ Rev.\ D~\textbf{91}, 116010 (2015);
  % doi:10.1103/PhysRevD.91.116010
  % arXiv:1502.08011~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevD.91.116010;%%
%
  R.~Lang, N.~Kaiser, and W.~Weise,
  % ``Shear viscosities from Kubo formalism in a large--$N_{c}$ Nambu--Jona--Lasinio model'',
  Eur.\ Phys.\ J.\ A~\textbf{51}, 127 (2015).
  % doi:10.1140/epja/i2015-15127-7
  % arXiv:1506.02459~[hep-ph].
  %%CITATION = doi:10.1140/epja/i2015-15127-7;%%


\bibitem{APT6}
  K.A.~Milton, I.L.~Solovtsov, and O.P.~Solovtsova,
  % ``The Bjorken sum rule in the analytic approach
  % to perturbative QCD,''
  Phys.\ Lett.\ B~\textbf{439}, 421 (1998);
  % [hep-ph/9809510].
  %%CITATION = HEP-PH/9809510;%%
%
  % K.A.~Milton, I.L.~Solovtsov, and O.P.~Solovtsova,
  % ``The Gross-Llewellyn Smith sum rule in the
  % analytic approach to perturbative QCD,''
  Phys.\ Rev.\ D~\textbf{60}, 016001 (1999);
  % [hep-ph/9809513].
  %%CITATION = HEP-PH/9809513;%%
%
  R.S.~Pasechnik, D.V.~Shirkov, and O.V.~Teryaev,
  % ``Bjorken Sum Rule and pQCD frontier on the move,''
  % Phys.\ Rev.\ D
  \textit{ibid.}~\textbf{78}, 071902 (2008);
  % [arXiv:0808.0066 [hep-ph]].
  %%CITATION = ARXIV:0808.0066;%%
%
  R.S.~Pasechnik, J.~Soffer, and O.V.~Teryaev,
  % ``Nucleon spin structure at low momentum transfers'',
  % Phys.\ Rev.\ D
  \textit{ibid.}~\textbf{82}, 076007 (2010).
  %[arXiv:1009.3355 [hep-ph]].
  %%CITATION = ARXIV:1009.3355;%%


\bibitem{APT7a}
  K.A.~Milton, I.L.~Solovtsov, and O.P.~Solovtsova,
  % ``Analytic perturbation theory and inclusive tau decay'',
  Phys.\ Lett.\ B~\textbf{415}, 104 (1997);
  % [hep-ph/9706409].
  %%CITATION = HEP-PH/9706409;%%
%
  K.A.~Milton, I.L.~Solovtsov, O.P.~Solovtsova, and V.I.~Yasnov,
  % ``Renormalization scheme and higher loop stability in hadronic
  % tau decay within analytic perturbation theory'',
  Eur.\ Phys.\ J.\ C~\textbf{14}, 495 (2000).
  % arXiv:hep-ph/0003030.
  %%CITATION = HEP-PH/0003030;%%


\bibitem{APT7b}
  G.~Ganbold,
  % ``Glueballs and Mesons: The Ground States'',
  Phys.\ Rev.\ D~\textbf{79}, 034034 (2009);
  % [arXiv:0903.0476 [hep-ph]].
  %%CITATION = ARXIV:0903.0476;%%
%
  % G.~Ganbold,
  % ``QCD Effective Coupling in the Infrared Region'',
  % Phys.\ Rev.\ D
  \textbf{81}, 094008 (2010);
  % [arXiv:1004.5280 [hep-ph]].
  %%CITATION = ARXIV:1004.5280;%%
%
  % G.~Ganbold,
  % ``Meson spectrum, glueball mass and QCD effective
  % coupling within infrared confinement'',
  % Proc.~of 10th International Conference on Quark Confinement
  % and the Hadron Spectrum (Munich, Germany, 2012);
  PoS~(Confinement~X), 065 (2012).
  %%CITATION = POSCI,ConfinementX,065;%%


\bibitem{APT8a}
  A.~Bakulev, K.~Passek--Kumericki, W.~Schroers, and N.~Stefanis,
  % ``Pion form-factor in QCD: From nonlocal condensates to NLO
  % analytic perturbation theory,''
  Phys.\ Rev.\ D~\textbf{70}, 033014 (2004);
  \textbf{70}, 079906(E) (2004);
  % arXiv:hep-ph/0405062.
  %%CITATION = HEP-PH/0405062;%%
%
  A.~Bakulev, A.~Pimikov, and N.~Stefanis,
  % ``QCD sum rules with nonlocal condensates and
  % the spacelike pion form factor,''
  % Phys.\ Rev.\ D
  \textit{ibid.}~\textbf{79}, 093010 (2009);
  % arXiv:0904.2304~[hep-ph].
  %%CITATION = ARXIV:0904.2304;%%
%
  N.~Stefanis,
  % ``Pion form--factor analysis using NLO analytic perturbation theory'',
  Nucl.\ Phys.\ B (Proc.\ Suppl.) \textbf{152}, 245 (2006).
  % arXiv:hep-ph/0410245.
  %%CITATION = HEP-PH/0410245;%%


\bibitem{APT8b}
  N.G.~Stefanis, W.~Schroers, and H.C.~Kim,
  % ``Pion form-factors with improved infrared factorization'',
  Phys.\ Lett.\ B~\textbf{449}, 299 (1999);
  % doi:10.1016/S0370-2693(99)00096-9
  % arXiv:hep-ph/9807298.
  %%CITATION = doi:10.1016/S0370-2693(99)00096-9;%%
%
  % N.G.~Stefanis, W.~Schroers, and H.C.~Kim,
  % ``Analytic coupling and Sudakov effects in exclusive processes:
  % Pion and gamma* gamma ---> pi0 form-factors'',
  Eur.\ Phys.\ J.\ C~\textbf{18}, 137 (2000);
  % doi:10.1007/s100520000510
  % arXiv:hep-ph/0005218.
  %%CITATION = doi:10.1007/s100520000510;%%
%
  A.I.~Karanikas and N.G.~Stefanis,
  % ``Analyticity and power corrections in hard scattering hadronic functions'',
  Phys.\ Lett.\ B~\textbf{504}, 225 (2001);
  \textbf{636}, 330(E) (2006);
  % doi:10.1016/S0370-2693(01)00297-0, 10.1016/j.physletb.2006.04.008
  % arXiv:hep-ph/0101031.
  %%CITATION = doi:10.1016/S0370-2693(01)00297-0, 10.1016/j.physletb.2006.04.008;%%
%
  A.P.~Bakulev, A.V.~Radyushkin, and N.G.~Stefanis,
  % ``Form-factors and QCD in space--like and time--like region'',
  Phys.\ Rev.\ D~\textbf{62}, 113001 (2000);
  % doi:10.1103/PhysRevD.62.113001
  % arXiv:hep-ph/0005085.
  %%CITATION = doi:10.1103/PhysRevD.62.113001;%%
%
  A.P.~Bakulev, A.I.~Karanikas, and N.G.~Stefanis,
  % ``Analyticity properties of three-point functions in QCD beyond leading order'',
  % Phys.\ Rev.\ D
  \textit{ibid.}~\textbf{72}, 074015 (2005).
  % doi:10.1103/PhysRevD.72.074015
  % arXiv:hep-ph/0504275.
  %%CITATION = doi:10.1103/PhysRevD.72.074015;%%


\bibitem{APT9}
  G.~Cvetic, R.~Kogerler, and C.~Valenzuela,
  % ``Reconciling the analytic QCD with the ITEP operator product expansion philosophy'',
  Phys.\ Rev.\ D~\textbf{82}, 114004 (2010);
  % [arXiv:1006.4199 [hep-ph]].
  %%CITATION = ARXIV:1006.4199;%%
%
  % G.~Cvetic, R.~Kogerler, and C.~Valenzuela,
  % ``Analytic QCD coupling with no power terms in UV regime'',
  J.\ Phys.\ G~\textbf{37}, 075001 (2010);
  % [arXiv:0912.2466 [hep-ph]].
  %%CITATION = ARXIV:0912.2466;%%
%
  G.~Cvetic and R.~Kogerler,
  % ``Applying generalized Pad\'e approximants in analytic QCD models'',
  Phys.\ Rev.\ D~\textbf{84}, 056005 (2011).
  % [arXiv:1107.2902 [hep-ph]].
  %%CITATION = ARXIV:1107.2902;%%


\bibitem{APT10}
  O.~Teryaev,
  % ``Analyticity and higher twists,''
  Nucl.\ Phys.\ B (Proc.\ Suppl.) \textbf{245}, 195 (2013);
  % arXiv:1309.1985~[hep-ph].
  %%CITATION = ARXIV:1309.1985;%%
%
  A.V.~Sidorov and O.P.~Solovtsova,
  % ``The QCD Analysis of $xF_3$ Structure Function Based on the Analytic Approach,''
  Nonlin.\ Phenom.\ Complex Syst.~\textbf{16}, 397 (2013);
  % arXiv:1312.3082~[hep-ph].
  %%CITATION = ARXIV:1312.3082;%%
%
  % A.V.~Sidorov and O.P.~Solovtsova,
  % ``The QCD analysis of the combined set for the $F_3$ structure
  % function data based on the analytic approach'',
  Mod.\ Phys.\ Lett.\ A~\textbf{29}, 1450194 (2014);
  % arXiv:1407.6858~[hep-ph].
  %%CITATION = ARXIV:1407.6858;%%
%
  % A.V.~Sidorov and O.P.~Solovtsova,
  % ``Non--singlet $Q^2$--evolution by inverse Mellin transform in
  % the analytic perturbation theory'',
  J.~Phys.\ Conf.\ Ser.~\textbf{678}, 012042 (2016);
  % doi:10.1088/1742-6596/678/1/012042
  %%CITATION = doi:10.1088/1742-6596/678/1/012042;%%
%
  % A.V.~Sidorov and O.P.~Solovtsova,
  % ``QCD analysis of the F$_{3}$ structure function based on inverse
  % Mellin transform in analytic perturbation theory'',
  Phys.\ Part.\ Nucl.\ Lett.~\textbf{14}, 1 (2017).
  % doi:10.1134/S154747711701023X
  %%CITATION = doi:10.1134/S154747711701023X;%%


\bibitem{12dAnQCD}
  C.~Contreras, G.~Cvetic, O.~Espinosa, and H.E.~Martinez,
  % ``Simple analytic QCD model with perturbative QCD behavior at high momenta'',
  Phys.\ Rev.\ D~\textbf{82}, 074005 (2010);
  % doi:10.1103/PhysRevD.82.074005
  % arXiv:1006.5050~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevD.82.074005;%%
%
  C.~Ayala, C.~Contreras, and G.~Cvetic,
  % ``Extended analytic QCD model with perturbative QCD behavior at high momenta'',
  % Phys.\ Rev.\ D
  \textit{ibid.}~\textbf{85}, 114043 (2012).
  % doi:10.1103/PhysRevD.85.114043
  % arXiv:1203.6897~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevD.85.114043;%%


\bibitem{PRD6264}
  A.V.~Nesterenko,
  % ``Quark--antiquark potential in the analytic approach to QCD'',
  Phys.\ Rev.\ D~\textbf{62}, 094028 (2000);
  % doi:10.1103/PhysRevD.62.094028
  % arXiv:hep-ph/9912351.
  %%CITATION = doi:10.1103/PhysRevD.62.094028;%%
%
  % A.V.~Nesterenko,
  % ``New analytic running coupling in spacelike and timelike regions'',
  % Phys.\ Rev.\ D~
  \textbf{64}, 116009 (2001).
  % doi:10.1103/PhysRevD.64.116009
  % arXiv:hep-ph/0102124.
  %%CITATION = doi:10.1103/PhysRevD.64.116009;%%


\bibitem{Review}
  A.V.~Nesterenko,
  % ``Analytic invariant charge in QCD'',
  Int.\ J.\ Mod.\ Phys.\ A~\textbf{18}, 5475 (2003);
  % doi:10.1142/S0217751X0301704X
  % arXiv:hep-ph/0308288.
  %%CITATION = doi:10.1142/S0217751X0301704X;%%
%
  % A.V.~Nesterenko,
  % ``Nonperturbative aspects of the QCD analytic invariant charge'',
  Nucl.\ Phys.\ B (Proc.\ Suppl.) \textbf{133}, 59 (2004).
  % doi:10.1016/j.nuclphysbps.2004.04.137
  % arXiv:hep-ph/0307283.
  %%CITATION = doi:10.1016/j.nuclphysbps.2004.04.137;%%


\bibitem{MPLA1516}
  A.V.~Nesterenko,
  % ``Investigation of a new analytic running coupling in~QCD'',
  Mod.\ Phys.\ Lett.\ A~\textbf{15}, 2401 (2000);
  % doi:10.1142/S0217732300003030
  % arXiv:hep-ph/0102203.
  %%CITATION = doi:10.1142/S0217732300003030;%%
%
  A.V.~Nesterenko and I.L.~Solovtsov,
  % ``New analytic running coupling in QCD: higher loop levels'',
  % Mod.\ Phys.\ Lett.\ A
  \textit{ibid.}~\textbf{16}, 2517 (2001).
  % doi:10.1142/S0217732301005989
  % arXiv:hep-ph/0111067.
  %%CITATION = doi:10.1142/S0217732301005989;%%


\bibitem{APTCSB}
  A.C.~Aguilar, A.V.~Nesterenko, and J.~Papavassiliou,
  % ``Infrared enhanced analytic coupling and chiral symmetry breaking in QCD'',
  J.~Phys.\ G~\textbf{31}, 997 (2005);
  % doi:10.1088/0954-3899/31/9/002
  % arXiv:hep-ph/0504195.
  %%CITATION = doi:10.1088/0954-3899/31/9/002;%%
%
  % A.C.~Aguilar, A.V.~Nesterenko, and J.~Papavassiliou,
  % ``Quark gap equation within the analytic approach to QCD'',
  Nucl.\ Phys.\ B (Proc.\ Suppl.) \textbf{164}, 300 (2007).
  % doi:10.1016/j.nuclphysbps.2006.11.071
  % arXiv:hep-ph/0510117.
  %%CITATION = doi:10.1016/j.nuclphysbps.2006.11.071;%%


\bibitem{RPert4L}
  P.A.~Baikov, K.G.~Chetyrkin, and J.H.~Kuhn,
  % ``Order $\alpha_{s}^{4}$ QCD corrections to~$Z$ and~$\tau$ decays'',
  Phys.\ Rev.\ Lett.~\textbf{101}, 012002 (2008);
  % doi:10.1103/PhysRevLett.101.012002
  % arXiv:0801.1821~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevLett.101.012002;%%
%
  % P.A.~Baikov, K.G.~Chetyrkin, and J.H.~Kuhn,
  % ``Adler function, Bjorken sum rule, and the Crewther relation
  % to order~$\alpha_{s}^{4}$ in a general gauge theory'',
  % Phys.\ Rev.\ Lett.~
  \textbf{104}, 132004 (2010);
  % doi:10.1103/PhysRevLett.104.132004
  % arXiv:1001.3606~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevLett.104.132004;%%
%
  P.A.~Baikov, K.G.~Chetyrkin, J.H.~Kuhn, and J.~Rittinger,
  % ``Adler function, sum rules and Crewther relation of order
  % $\co(\alpha_{s}^{4})$: the singlet case'',
  Phys.\ Lett.\ B~\textbf{714}, 62 (2012).
  % doi:10.1016/j.physletb.2012.06.052
  % arXiv:1206.1288~[hep-ph].
  %%CITATION = doi:10.1016/j.physletb.2012.06.052;%%


\bibitem{Beta5L}
  P.A.~Baikov, K.G.~Chetyrkin, and J.H.~Kuhn,
  % ``Five--loop running of the QCD coupling constant'',
  Phys.\ Rev.\ Lett.~\textbf{118}, 082002 (2017);
  % doi:10.1103/PhysRevLett.118.082002
  % arXiv:1606.08659~[hep-ph].
  %%CITATION = doi:10.1103/PhysRevLett.118.082002;%%
%
  F.~Herzog, B.~Ruijl, T.~Ueda, J.A.M.~Vermaseren, and A.~Vogt,
  % ``The five-loop beta function of Yang--Mills theory with fermions'',
  JHEP~\textbf{1702}, 090 (2017).
  % doi:10.1007/JHEP02(2017)090
  % arXiv:1701.01404~[hep-ph].
  %%CITATION = doi:10.1007/JHEP02(2017)090;%%


\bibitem{CPC}
  A.V.~Nesterenko and C.~Simolo,
  % ``QCDMAPT: Program package for analytic approach to QCD'',
  Comput.\ Phys.\ Commun.~\textbf{181}, 1769 (2010);
  % doi:10.1016/j.cpc.2010.06.040
  % arXiv:1001.0901~[hep-ph].
  %%CITATION = doi:10.1016/j.cpc.2010.06.040;%%
%
  % A.V.~Nesterenko and C.~Simolo,
  % ``QCDMAPT\underline{~~}F:~Fortran version of QCDMAPT package'',
  % Comput.\ Phys.\ Commun.~
  \textbf{182}, 2303 (2011).
  % doi:10.1016/j.cpc.2011.05.020
  % arXiv:1107.1045~[hep-ph].
  %%CITATION = doi:10.1016/j.cpc.2011.05.020;%%


\bibitem{Pi2termsHO}
  A.V.~Nesterenko and S.A.~Popov,
  % ``$R$--ratio of electron--positron annihilation into hadrons:
  % higher--order~$\pi^2$--terms'',
  Nucl.\ Part.\ Phys.\ Proc.~\textbf{282}, 158 (2017).
  % doi:10.1016/j.nuclphysbps.2016.12.030
  % arXiv:1608.04205~[hep-ph].
  %%CITATION = doi:10.1016/j.nuclphysbps.2016.12.030;%%


\bibitem{RpertHO}
  A.V.~Nesterenko,
  % ``Electron--positron annihilation into hadrons at the higher--loop levels'',
  Eur.\ Phys.\ J.\ C~\textbf{77}, 844 (2017).
  % doi:10.1140/epjc/s10052-017-5405-5
  % arXiv:1707.00668~[hep-ph].
  %%CITATION = ARXIV:1707.00668;%%


\bibitem{Pi2Terms1}
  J.D.~Bjorken,
  % ``Two~topics in Quantum Chromodynamics'',
  report SLAC--PUB--5103 (1989).
  %%CITATION = SLAC-PUB-5103;%%


\bibitem{RPert5LEstim1}
  A.L.~Kataev and V.V.~Starshenko,
  % ``Estimates of the higher order QCD corrections to~$R(s)$, $R_{\tau}$,
  % and deep inelastic scattering sum rules'',
  Mod.\ Phys.\ Lett.\ A~\textbf{10}, 235 (1995).
  % doi:10.1142/S0217732395000272
  % arXiv:hep-ph/9502348.
  %%CITATION = doi:10.1142/S0217732395000272;%%


\bibitem{ProsperiAlpha}
  G.M.~Prosperi, M.~Raciti, and C.~Simolo,
  % ``On~the running coupling constant in~QCD'',
  Prog.\ Part.\ Nucl.\ Phys.~\textbf{58}, 387 (2007).
  % doi:10.1016/j.ppnp.2006.09.001
  % arXiv:hep-ph/0607209.
  %%CITATION = doi:10.1016/j.ppnp.2006.09.001;%%


\end{thebibliography}

\documentclass[a4paper]{article}

\usepackage{18lomcon}        
\usepackage{cite}            
\usepackage{epsfig}          
\usepackage{amsmath,amssymb}

\newcommand{\ind}[2]{^{#1}_{\text{#2}}}
\def\epem{e^{+}e^{-}}
\def\nf{n_{f}}

\begin{document}

\title{CALCULATION OF THE $R$--RATIO OF {\large$\epem\!\to\,$}HADRONS \\[1mm]
AT THE HIGHER--LOOP LEVELS}

\author{A.V.~Nesterenko \email{nesterav@theor.jinr.ru,~nesterav@gmail.com}}

\affiliation{BLTPh JINR, Dubna, 141980, Russian Federation}

\maketitle

\begin{abstract}
The calculation of the $R$--ratio of electron--positron annihilation into
hadrons is discussed. The method, which enables one to properly account
for all the effects due to continuation of the spacelike perturbative
results into the timelike domain at an arbitrary loop level, is
delineated.
\end{abstract}

The theoretical description of a variety of the strong interaction
processes is inherently based on the hadronic vacuum polarization
function~$\Pi(q^2)$, which is defined as the scalar part of the hadronic
vacuum polarization tensor
\begin{equation}
\label{P_Def}
\Pi_{\mu\nu}(q^2) = i \!\int\! d^4 x\,e^{i q x} \bigl\langle 0 \bigl|\,
T\left\{J_{\mu}(x)\, J_{\nu}(0)\right\} \bigr| 0 \bigr\rangle =
\frac{i}{12\pi^2} (q_{\mu}q_{\nu} - g_{\mu\nu}q^2) \Pi(q^2),
\end{equation}
the related Adler function~\cite{Adler}
\begin{equation}
\label{Adler_Def}
D(Q^2) = - \frac{d\, \Pi(-Q^2)}{d \ln Q^2},
\end{equation}
and the function~$R(s)$
\begin{equation}
\label{R_Def}
R(s) = \frac{1}{2 \pi i} \lim_{\varepsilon \to 0_{+}}
\Bigl[\Pi(s + i \varepsilon) - \Pi(s - i \varepsilon)\Bigr] =
\frac{\sigma(\epem \to \text{hadrons}; s)}
{\sigma(\epem \to \mu^{+}\mu^{-}; s)},
\end{equation}
which is identified with the so--called $R$--ratio of electron--positron
annihilation into hadrons. The functions~$\Pi(q^2)$~(\ref{P_Def})
and~$D(Q^2)$~(\ref{Adler_Def}), being the functions of the spacelike
kinematic variable~$Q^2 = -q^2 > 0$, can be directly accessed within
QCD~perturbation theory, whereas the $R$--ratio~(\ref{R_Def}), being the
function of the timelike kinematic variable~$s = q^2 > 0$, can be
described only by making use of the relevant dispersion relations.
Specifically, the relation, which expresses the $R$--ratio~(\ref{R_Def})
in terms of the theoretically calculable Adler function and provides a
native way to properly account for the effects due to continuation of the
spacelike perturbative results into the timelike domain, can be obtained
by integrating Eq.~(\ref{Adler_Def}) in finite limits, that
yields~\cite{Rad82, KP82}
\begin{equation}
\label{R_Disp2}
R(s) =  \frac{1}{2 \pi i} \lim_{\varepsilon \to 0_{+}}
\int_{s + i \varepsilon}^{s - i \varepsilon}
D(-\zeta)\,\frac{d \zeta}{\zeta}.
\end{equation}
The integration contour in this equation lies in the region of analyticity
of the integrand in the complex $\zeta$--plane.

It is necessary to outline that the dispersion relations, which express
the functions~$\Pi(q^2)$, $R(s)$, and~$D(Q^2)$ in terms of each other,
rely only on the kinematics of the process on hand and involve neither
model--dependent phenomenological assumptions nor additional
approximations. In~turn, such relations impose a number of strict physical
intrinsically nonperturbative constraints on the functions~$\Pi(q^2)$,
$R(s)$, and~$D(Q^2)$, that should certainly be accounted for when one
comes out of the limits of applicability of the QCD perturbation theory.
It is worthwhile to note also that these nonperturbative restrictions have
been merged with corresponding perturbative input in the framework of
dispersively improved perturbation theory~(DPT)~\cite{Book, DPT1, DPT2}
(its preliminary formulation was discussed in Ref.~\cite{DPTPrelim}). In
particular, the DPT enables one to overcome some inherent difficulties of
the QCD perturbation theory and extend its applicability range towards the
infrared domain, see book~\cite{Book} and references therein.

This study is primarily focused on the theoretical description of the
$R$--ratio of electron--positron annihilation into hadrons~(\ref{R_Def})
at moderate and high energies, so that the nonperturbative aspects of the
strong interactions will be disregarded hereinafter. For this purpose the
effects due to the masses of the involved particles can be safely
neglected (the impact of such effects on the low--energy behavior of the
functions on hand was discussed in, e.g., Refs.~\cite{Book, DPT1, DPT2,
DPT3}). In the massless limit the relation~(\ref{R_Disp2}) can be
represented as
\begin{equation}
\label{RProp}
R^{(\ell)}(s) = 1 + r^{(\ell)}(s),
\quad
r^{(\ell)}(s) =
\int_{s}^{\infty}\rho^{(\ell)}(\sigma)\,
\frac{d \sigma}{\sigma},
\end{equation}
where
\begin{equation}
\label{RhoDef}
\rho^{(\ell)}(\sigma) =
\frac{1}{2 \pi i} \lim_{\varepsilon \to 0_{+}}
\Bigl[d^{(\ell)}(-\sigma - i \varepsilon) -
d^{(\ell)}(-\sigma + i \varepsilon)\Bigr]
\end{equation}
stands for the spectral function and~$d^{(\ell)}(Q^2)$ denotes the
$\ell$--loop strong correction to the Adler function~(\ref{Adler_Def}).
As~mentioned above, only perturbative contributions will be retained in
Eq.~(\ref{RhoDef}) in what follows, that makes Eq.~(\ref{RProp}) identical
to that of the ``Analytic perturbation theory''~\cite{APT0} (for some of
its applications see Refs.~\cite{APT1, APT2, APT3, APT4, APT5, APT6,
APT7a, APT7b, APT8a, APT8b, APT9, APT10, 12dAnQCD}). A discussion of the
nonperturbative terms in the spectral density~$\rho^{(\ell)}(\sigma)$ can
be found in, e.g., Refs.~\cite{PRD6264, Review, MPLA1516, APTCSB}.

The perturbative expression for the Adler function~(\ref{Adler_Def}) takes
the form of the power series in the so--called QCD couplant
$a\ind{(\ell)}{s}(Q^2) = \alpha\ind{(\ell)}{s}(Q^2)\, \beta_{0}/(4\pi)$
\begin{equation}
\label{AdlerPert}
D^{(\ell)}_{{\rm pert}}(Q^2) = 1 + d^{(\ell)}_{{\rm pert}}(Q^2),
\quad
d^{(\ell)}_{{\rm pert}}(Q^2) = \sum_{j=1}^{\ell} d_{j}
\left[a\ind{(\ell)}{s}(Q^2)\right]^{j}.
\end{equation}
Here~$\ell$ specifies the loop level, $d_1 = 4/\beta_0$, $\beta_0 = 11 -
2\nf/3$, $\nf$~is the number of active flavors, the common prefactor
$N_{\text{c}}\sum_{f=1}^{\nf} Q_{f}^{2}$ is omitted throughout,
$N_{\text{c}}=3$ denotes the number of colors, and $Q_{f}$~stands for the
electric charge of $f$--th quark. The QCD couplant entering
Eq.~(\ref{AdlerPert}) can be represented~as
\begin{equation}
\label{AItGen}
a\ind{(\ell)}{s}(Q^2) =
\sum_{n=1}^{\ell}\sum_{m=0}^{n-1} b^{m}_{n}\,
\frac{\ln^{m}(\ln z)}{\ln^n z},
\end{equation}
where $z=Q^2/\Lambda^2$, $b^{m}_{n}$ is the combination of the
$\beta$~function perturbative expansion coefficients ($b^{0}_{1}=1$,
$b^{0}_{2}=0$, $b^{1}_{2}=-\beta_{1}/\beta^{2}_{0}$,~etc.), and~$\Lambda$
denotes the QCD scale parameter. The Adler function perturbative expansion
coefficients~$d_{j}$ were calculated up to the four--loop
level~\cite{RPert4L}, whereas the $\beta$~function perturbative expansion
coefficients~$\beta_{j}$ are available up to the five--loop
level~\cite{Beta5L}.

Since the calculation of the spectral
function~$\rho^{(\ell)}(\sigma)$~(\ref{RhoDef}) becomes rather cumbrous
beyond the one--loop level (the explicit expressions
for~$\rho^{(\ell)}(\sigma)$ at first four loop levels can be found in
Ref.~\cite{CPC}), one commonly re--expands the strong
correction~$r^{(\ell)}(s)$~(\ref{RProp}) at high energies, that eventually
leads to~\cite{Pi2termsHO, RpertHO}
\begin{align}
\label{Pi2TermsGen}
r^{(\ell)}(s) & =
\sum_{j=1}^{\ell} d_{j} \left[a^{(\ell)}_{{\rm s}}(|s|)\right]^{j} -
\sum_{j=1}^{\ell} d_{j}
\sum_{n=1}^{\infty}\! \frac{(-1)^{n+1}}{(2n+1)!}\pi^{2n} \times
\nonumber \\ & \hspace{-7mm} \times
\!\sum_{k_{1}=0}^{\ell-1}
\!\ldots\!
\sum_{k_{2n}=0}^{\ell-1}\!
\left(\prod_{p=1}^{2n}B_{k_{p}}\!\right)\!
\left[\prod_{t=0}^{2n-1}\!
\Bigl(j+t+k_{1}+k_{2}+\ldots+k_{t}\Bigr)\!\right]\! \times
\nonumber \\ & \hspace{-7mm} \times\!
\Bigl[a^{(\ell)}_{{\rm s}}(|s|)\Bigr]^{j+2n+k_{1}+k_{2}+\ldots+k_{2n}},
\qquad \sqrt{s}/\Lambda>\exp(\pi/2).
\end{align}
It is necessary to emphasize that the re--expansion~(\ref{Pi2TermsGen}) is
valid only for~$\sqrt{s}/\Lambda > \exp(\pi/2) \simeq 4.81$, and it
converges rather slowly when the energy scale approaches this value. If
the number of terms retained on the right--hand side of
Eq.~(\ref{Pi2TermsGen}) is large enough, then it can provide quite
accurate approximation of the strong correction to
the~$R$--ratio~(\ref{RProp}). However, one usually truncates the
re--expansion~(\ref{Pi2TermsGen}) at the order~$\ell$, thereby neglecting
all the higher--order $\pi^2$--terms (though, the latter may not
necessarily be negligible, see Ref.~\cite{RpertHO}), that results in the
expression commonly employed in the practical applications:
\begin{equation}
\label{RAppr}
R^{(\ell)}_{{\rm appr}}(s) = 1 + r^{(\ell)}_{{\rm appr}}(s),
\quad
r^{(\ell)}_{{\rm appr}}(s) =\!
\sum_{j=1}^{\ell} r_{j} \!\left[a^{(\ell)}_{{\rm s}}(|s|)\right]^{j}\!,
\quad
r_{j} = d_{j} - \delta_{j}.
\end{equation}
Here $d_j$ denote the Adler function perturbative expansion
coefficients~(\ref{AdlerPert}) and~$\delta_{j}$ embody the contributions
of relevant~$\pi^2$--terms~(\ref{Pi2TermsGen}), see Refs.~\cite{Pi2Terms1,
RPert5LEstim1, ProsperiAlpha, Book, Pi2termsHO, RpertHO}.

At the same time, the explicit expression for the perturbative spectral
function entering Eq.~(\ref{RProp}) can be calculated at an arbitrary loop
level (it is assumed that the involved perturbative coefficients~$d_j$
and~$\beta_j$ are known) by making use of the method developed in
Ref.~\cite{RpertHO}, namely
\begin{align}
\label{RhoPertHO}
\rho^{(\ell)}(\sigma) & =
\sum_{j=1}^{\ell} d_{j} \sum_{k=0}^{K(j)}
\binom{j}{2k+1} (-1)^{k}\, \pi^{2k} \times
\nonumber \\
& \times
\Biggl[\sum_{n=1}^{\ell}\sum_{m=0}^{n-1} b^{m}_{n}\, u_{n}^{m}(\sigma)\Biggr]^{j-2k-1}
\Biggl[\sum_{n=1}^{\ell}\sum_{m=0}^{n-1} b^{m}_{n}\, v_{n}^{m}(\sigma)\Biggr]^{2k+1}.
\end{align}
In this equation
\begin{align}
\label{Unm}
u_{n}^{m}(\sigma) & =
\begin{cases}
u_{n}^{0}(\sigma),& \text{if $\, m=0$},\\
u_{n}^{0}(\sigma) u_{0}^{m}(\sigma) -
\pi^2 v_{n}^{0}(\sigma) v_{0}^{m}(\sigma),\quad& \text{if $\, m \ge 1$},
\end{cases}
\\
\label{Vnm}
v_{n}^{m}(\sigma) & =
\begin{cases}
v_{n}^{0}(\sigma),& \text{if $\, m=0$},\\
v_{n}^{0}(\sigma) u_{0}^{m}(\sigma) +
u_{n}^{0}(\sigma) v_{0}^{m}(\sigma),\quad\quad\,& \text{if $\, m \ge 1$},
\end{cases}
\\
\label{Un0Def}
u_{n}^{0}(\sigma) & = \frac{1}{(y^2 + \pi^2)^{n}}
\sum_{k=0}^{K(n+1)} \binom{n}{2k} (-1)^{k} \pi^{2k} y^{n-2k},
\\
\label{Vn0Def}
v_{n}^{0}(\sigma) & = \frac{1}{(y^2 + \pi^2)^{n}}
\sum_{k=0}^{K(n)} \binom{n}{2k+1} (-1)^{k} \pi^{2k} y^{n-2k-1},
\\
\label{U0mDef}
u_{0}^{m}(\sigma) & = \sum\limits_{k=0}^{K(m+1)}\binom{m}{2k}
(-1)^{k}\pi^{2k} \Bigl[L_{1}(y)\Bigr]^{m-2k}\, \Bigl[L_{2}(y)\Bigr]^{2k},
\\
\label{V0mDef}
v_{0}^{m}(\sigma) & = \sum\limits_{k=0}^{K(m)}\binom{m}{2k+1}
(-1)^{k+1}\pi^{2k} \Bigl[L_{1}(y)\Bigr]^{m-2k-1}\, \Bigl[L_{2}(y)\Bigr]^{2k+1},
\end{align}
\begin{equation}
\label{L12Def}
L_{1}(y) \!=\! \ln\!\sqrt{y^{2}+\pi^{2}},
\quad
L_{2}(y) \!=\! \frac{1}{2} - \frac{1}{\pi}\arctan\!\left(\frac{y}{\pi}\right),
\end{equation}
$K(j) = [(j-2)+(j\;\text{mod}\;2)]/2$, $y=\ln(\sigma/\Lambda^2)$, and~$n
\ge 1$ is assumed. In~turn, Eq.~(\ref{RhoPertHO}) enables one to properly
account for the effects due to continuation of the spacelike perturbative
results into the timelike domain at an arbitrary loop level, that plays a
valuable role in the studies of a variety of the strong interaction
processes, see paper~\cite{RpertHO} and references therein for the details.

\end{document}